Jetson Spacecraft Corp. shows the following information on its 2015 income statement: sales = $318944; costs = $153087; other expenses = $6405; depreciation expense = $19056; interest expense = $14042; taxes = $18299; dividends = $11761. In addition, you’re told that the firm issued $5054 in new equity during 2015 and redeemed $5683 in outstanding long-term debt. If net fixed assets increased by $21297 during the year, what was the addition to NWC?
Answer #1


Cash Flow from Operations = EBIT - Taxes + Depreciation = (Sales - Costs - Other Expenses - Depreciation) -Taxes + Depreciation = (318,944 - 153087 - 6405 -19056) - 18299 + 19056 = 141,153

Cash Flow from Assets = Cash Flow from Creditors + Cash Flow Equity holders = (Interest - New Debt ) +(Dividends - New Equity) = (14042 - 5683) +(11761 - 5054) = 15066

Cash Flow from operations = Cash Flow from Assets - Increase in Fixed Assets - Increase in NWC
141153 = 15066 - 21297 - Increase in NWC
Increase in NWC = 15066 - 21297 -141153 = -147,384

Decrease in NWC = 147,384

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well